OpenStudy (anonymous):

Estimate 78.65 - 9.5. Explain your method.

OpenStudy (anonymous):

stupid question

OpenStudy (hitaro9):

Rounding is pretty central to estimations. When a number's decimal portion is .5 or higher, then it is closer to the higher number, similarly, if it is lower than .5, then it is closer to the lower number. So you can round 78.65 to 79, and round 9.5 to 10 79-10 = 69

OpenStudy (anonymous):

Round each term to the nearest unit, 1, and then subtract. 79 - 10 = 69 78.65 - 9.5 = 69.15

OpenStudy (anonymous):

Sorry. @Hitaro9 You posted first.